I'm a school student. Any help would be greatly appreciated.

1)Firm under perfect competitive market. Equilibrium quantity of output is 200 units, Average variable cost is RS.400, Average revenue is RS.10. find the production surplus.

2)Rice is imported at RS.200 per kilo and sold at RS.100 per kilo. Demand equation is expressed as

P=600-0.5Qd. How much is the dead weight loss.

3)Household consumption expenditure of a household which earns an annual disposable income of RS.60000 is RS.45000. this indicates

Marginal propensity to consume 0.25

Marginal propensity to consume is 0.75

Average propensity to consume is 0.25

Average propensity to consume is 0.75

Marginal propensity to savings is 0.75

Which one.

4)Which of the following would cause a reduction in equilibrium national income,

Reduction is government taxes and increase in government expenditure

Decrease in marginal propensity to consume from 0.8 to 0.75

Reduction in marginal propensity savings from 0.25 to 0.20

Increase in government expenditure and transfer payments







1
Expert's answer
2021-01-07T10:53:26-0500

1) If equilibrium Q is 200 units, AVC is RS.400, AR is RS.10, then the production surplus is: "PS = 0.5\u00d710\u00d7200 = RS.1,000."

2) If demand equation is expressed as P = 600 - 0.5Qd or Qd = 1,200 - 2P, then the dead weight loss is:

"DWL = 0.5\u00d7(1,000 - 800)\u00d7(200 - 100) = RS.10,000."

3) If a household earns an annual disposable income of RS.60000 and spends RS.45000, then this indicates, that marginal propensity to consume is

"mpc = 45,000\/60,000 = 0.75."

So, the correct answer is b.

4) Reduction in marginal propensity savings from 0.25 to 0.20 would cause a reduction in equilibrium national income.

So, the correct answer is c.